A style option for rotated objects in L^{A}TEX^{∗}

Leonor Barroca

1995/04/07

Contents

1 History 1

2 Usage 2

3 Setup 2

List of Figures

List of Tables

Abstract

This article documents a L^{A}TEX package, ‘rotating.sty’, which perform all the different sorts of rotation one might like, including complete figures.

1 History

Version 2.0 is a complete re-write, with most of the work now being done by the L^{A}TEX 2ε graphics package.

Version 2.1 provides a ‘clockwise’ option to reinstate the behaviour described in the ‘L^{A}TEX Companion’

Version 2.2 just intercepts the standard float macros instead of copying and changing the. The ‘twoside’ option is obeyed.

Version 2.5 corrects problems in sideways figures.

Version 2.6 is a rewrite of the sideways floats via Frank Mittelbach (to whom many thanks for lookingat the mangy code).

Version 2.7 is checked for L^{A}TEX of December 94, and adds the option of twoside behaviour independent of the general twoside.

Version 2.8 cleans up some mistakes pointed out by Harald Axel Sommerfeldt. Version 2.9 cleans up some (more) mistakes pointed out by Harald Axel Sommerfeldt.

∗This file has version number 2.9, last revised 1995/04/07.

2 Usage

This style option provides three L^{A}TEX environments: sideways prints the contents turned through 90 degrees counterclockwise turn prints the contents turned through an arbitrary angle rotate prints the contents turned through an arbitrary angle but does not leave

any space for the result A full set of examples are given in the file examples.tex Now we present the documented code.

3 Setup